OpenStudy (anonymous):

Find the equation of the line parallel to 3x+4y+12=0 and passing at a distance 6 from the origin

OpenStudy (bradely):

Calculate the slope of the line 3x+4y+12=0 m = -a/b when equation is in the form ax+by+c=0 m= -3/4 Slope of the parallel lines are equal y= (-3/4)x+c 3x+4y-4c=0 Distance from the origin is 6 |3(0)+4(0)-4c|/sqrt(3^2+4^2) =6 4c/5=6 4c=30 c=30/4 =15/2 y= (-3/4)x+15/2 Source:www.mathskey.com